Please help me with this problem. Perform the requested operation or operations
f(x)=4x+11,g(x)==5x-1 find (f g)(x). WHere it says Find which is underlined there is a small "o" in between the f and G Thank you

Answer by malaydassharma(59)


f(x)= 4x + 11
g(x)= 5 x - 1
(fg)(x)=f[g(x)]
=f[5x-1] as g(x)= 5 x - 1

= 4(5x-1) +11 as f(x)= 4x + 11 and here x=(5x-1)

=20x +7